(1 point)

An urn has 3 red, 5 white, and 3 blue balls in an urn and two are drawn one after another without replacement.

What is the probability that the two balls are both red given that they are the same color?
Answer #1

P(Both red) = 3/11 * 2 / 10 = 3/55

P(Same color ) = P(both red) + P(both white) + P(both blue)

= 3/11 * 2/10 + 5/11 * 4/10 + 3/11 * 2/10

= 16/55

P(Both red | Both same color) = P(Both red) / P(Both same color)

= (3/55) / (16/55)

= 3 / 16 ( = 0.1875)